# font=, innertopmargin=2pt, innerbottommargin=2pt, innerleftmargin=8pt, innerrightmargin=8pt, frametitleaboveskip=2pt, frametitlebelowskip=1pt, frametitlealignment=, middlelinecolor=blue

### v2

@@ 4.2 Main Results and Analysis
The results of these evaluations are detailed in the metrics reported in Table 1. Through the comparison, it can be observed that although LLMob may not excel in replicating spatial features (SD) precisely, it demonstrates superior performance in handling temporal aspects (DI). When considering spatial-temporal features (DARD and STVD), LLMob’s performance is also competitive. In particular, LLMob achieves the best performance on DI and DARD for all the three settings, and is the runner-up on STVD. Baselines like DeepMove and TrajGAIL perform the best on SD and STVD, respectively, but become much less competitive when evaluated in other aspects. We suggest that the pronounced advantage of LLMob in terms of DARD (roughly 1/2 to 1/3 JSD compared to the best of baselines) can be attributed to the LLM agent’s tendency to accurately replicate the motivation behind individual activity behaviors. For instance, an agent may recognize patterns like a person’s habits to have breakfast in the morning, without being restricted to a specific restaurant. This phenomenon highlights the enhanced semantic understanding capabilities of the LLM agent.
@@ 5 Conclusion
Contributions. This study is believed to be the first personal mobility simulation empowered by LLM agents on real-world data. Our innovative framework leverages activity patterns and motivations to direct LLM agents in emulating urban residents, facilitating the generation of interpretable and effective individual activity trajectory. Extensive experimental studies based on real-world data are conducted to validate the proposed framework and demonstrate the promising capabilities of LLM agents to improve urban mobility analysis.
Social Impacts. Leveraging artificial intelligence to enhance societal benefits is increasingly promising, especially with the advent of high-capacity models such as LLMs. This study explores one of the potential avenues for applications using LLMs as reliable agents to simulate specific scenarios to assess the effects of external factors, such as pandemic and government policies. The introduced framework offers a flexible approach to enhance the reliability of LLMs in simulating urban mobility.
Limitations. In this study, we considered individual agents’ activities without interactions between them. We will target the multi-agent case to model interactions (e.g., people may follow others in a friends’ or family’s activity) as future work. Due to the difficulty in collecting high-quality personal mobility data (most datasets are incomplete in showing one’s daily activity), only one dataset is used in the experimental evaluation. Moreover, only GPT-3.5 was evaluated for the sake of cost-efficiency, whereas more advanced LLMs may yield better generation performance.
@@ Appendix C Experimental Setup
C.1 Environment
@@
C.2 Personal Activity Dataset
We use a personal activity trajectory dataset from Tokyo which was obtained through Twitter and Foursquare APIs, covering the data from January 2019 to December 2022. We randomly choose 100 users to model their individual activity trajectory at a 10-minute interval according to the number of available trajectories. Samples are shown in Table 2.
Table 2: Samples of personal activity data.
UserID Latitude Longitude Location Name Category Time
44673 35.008 139.015 Convenience Store Shop & Service 2019-12-17 8:00
44673 35.009 139.018 Ramen Restaurant Food & Service 2019-12-17 8:30
44673 35.004 139.060 Italian Restaurant Food & Service 2019-12-17 11:20
44673 35.009 139.085 Farmers Market Shop & Service 2019-12-17 14:20
44673 35.005 139.086 Soba Restaurant Food & Service 2019-12-17 18:00

### v3

@@ 4.2 Main Results and Analysis
The results of these evaluations are detailed in the metrics reported in Table 2. Through the comparison, it can be observed that although LLMob may not excel in replicating spatial features (SD) precisely, it demonstrates superior performance in handling temporal aspects (DI). When considering spatial-temporal features (DARD and STVD), LLMob’s performance is also competitive. In particular, LLMob achieves the best performance on DI and DARD for all three settings and is the runner-up on STVD. Baselines like DeepMove and TrajGAIL perform the best on SD and STVD, respectively, but become much less competitive when evaluated in other aspects. We suggest that the pronounced advantage of LLMob in terms of DARD (roughly 1/2 to 1/3 JSD compared to the best of baselines) can be attributed to the LLM agent’s tendency to accurately replicate the motivation behind individual activity behaviors. For instance, an agent may recognize patterns like a person’s habits to have breakfast in the morning, without being restricted to a specific restaurant. This phenomenon highlights the enhanced semantic understanding capabilities of the LLM agent.
@@ 5 Conclusion
Contributions. This study is believed to be the first personal mobility simulation empowered by LLM agents on real-world data. Our innovative framework leverages activity patterns and motivations to direct LLM agents in emulating urban residents, facilitating the generation of interpretable and effective individual activity trajectories. Extensive experimental studies based on real-world data are conducted to validate the proposed framework and demonstrate the promising capabilities of LLM agents to improve urban mobility analysis.
Social Impacts. Leveraging artificial intelligence to enhance societal benefits is increasingly promising, especially with the advent of high-capacity models such as LLMs. This study explores one of the potential avenues for applications using LLMs as reliable agents to simulate specific scenarios to assess the effects of external factors, such as pandemics and government policies. The introduced framework offers a flexible approach to enhance the reliability of LLMs in simulating urban mobility.
Limitations. In this study, we focused on modeling the activities of individual agents without considering interactions between them. As future work, we aim to extend this to a multi-agent scenario to capture interactions (e.g., where individuals may follow the activities of friends or family members). Given the challenges in collecting high-quality personal mobility data—many datasets lack completeness in capturing daily activities—we limited our comprehensive experimental evaluation to a single dataset. Furthermore, due to cost-efficiency considerations, only GPT-3.5 was fully evaluated. An additional analysis in a different city is provided in Appendix D.3, and Appendix D.4 includes supplementary evaluations using other LLMs.
Acknowledgements
This work is supported by JSPS KAKENHI JP22H03903, JP23H03406, JP23K17456, JP24K02996, and JST CREST JPMJCR22M2.
@@ Appendix C Experimental Setup
C.1 Data processing
All the data is obtained through the Twitter and Foursquare APIs and is already anonymized to remove any personally identifiable information before analysis. The detailed process is as follows:
  1. 1.
    Filtering Incomplete Data
    Users with missing check-ins for a specific year were filtered out.
  2. 2.
    Excluding Non-Japan Check-ins
    Check-ins that occurred outside of Japan were removed.
  3. 3.
    Inferring Prefecture from GPS Coordinates
    Prefectures were inferred based on the latitude and longitude data of check-ins.
  4. 4.
    Assigning Prefecture
    Users were assigned to a prefecture based on their primary check-in location; for example, users whose top check-in location is Tokyo were categorized as belonging to Tokyo.
  5. 5.
    Removing Sudden-Move Check-ins
    Check-ins showing abrupt, unrealistic location changes, such as from Tokyo to the United States within a short time frame, were deleted to remove data drift, following the criteria proposed by [40].
  6. 6.
    Anonymizing Data
    Real user IDs and geographic location names were anonymized. Only category information of geographic locations was kept, and latitude and longitude coordinates were converted into IDs before being input into the model.
C.2 Environment
@@ D.2 Examples of Retrieved Motivations and Corresponding Generated Trajectories
D.3 Experiment on Osaka Data
We conducted an experiment based on the data collected in Osaka, Japan. We generated 537 trajectories based on the 2102 daily activity trajectories from 30 persons. The results are reported as follows, where LLMob-L/E are ours and DiffTraj and TrajGAIL are the best-performing baseline methods.
Table 5: Comparison of models based on various metrics on Osaka data
Model SD SI DARD STVD
LLMob-L 0.035 0.021 0.141 0.391
LLMob-E 0.030 0.018 0.121 0.380
DiffTraj 0.080 0.177 0.406 0.691
TrajGAIL 0.281 0.063 0.525 0.483
D.4 Experiment on different LLMs
We conducted experiments for setting (1) using different LLMs (GPT-4o-mini and Llama 3-8B). The results are reported as follows:
Table 6: Results of experiments using different LLMs
Model SD SI DARD STVD
LLMob-L (GPT-3.5-turbo) 0.049 0.054 0.136 0.570
LLMob-L (GPT-4o-mini) 0.049 0.055 0.141 0.577
LLMob-L (Llama 3-8B) 0.054 0.063 0.119 0.566
LLMob-E (GPT-3.5-turbo) 0.053 0.046 0.125 0.559
LLMob-E (GPT-4o-mini) 0.041 0.053 0.211 0.531
LLMob-E (Llama 3-8B) 0.054 0.059 0.122 0.561
We observe competitive performance of our framework when other LLMs are used. In particular, GPT-4o-mini is the best in terms of the spatial metric (SD); GPT-3.5-turbo is the best in terms of the temporal metric (SI). Llama 3-8B is overall the best when spatial and temporal factors are evaluated together (DARD and STVD). Such results demonstrate the robustness of our framework across different LLMs.